# ETH-monotonicity and the black hole singularity

### 11 Mar 2026 v3

@@ 5 Quantum fluctuations
The Taylor expansions of $S(E_{n}+\omega)$ and $S(E_{n}+\omega/2)$ in (3.7) have higher derivative terms $S^{\prime\prime}(\bar{E})$, $S^{\prime\prime\prime}(\bar{E})$, etc. $S^{\prime\prime}(\bar{E})$ gives a correction which is of the order $O(G^{1}_{N})$. $S^{\prime\prime\prime}(\bar{E})$ and higher derivative terms are suppressed by higher positive powers of $G_{N}$. So, we will concentrate on the contribution from $S^{\prime\prime}(\bar{E})=-\beta^{2}/C$, where $C$ is the total heat capacity Murthy:2019fgs . It is due to quantum fluctuations which mimic thermal fluctuations Srednicki_1999 . So far, we have been ignoring this term. We can Taylor expand the exponential of this term due to the presence of $e^{\beta\omega/2}$ where $\beta\omega/2\sim O(1)$. The extra contribution to the change in energy due to the quantum fluctuations is given by
@@
It can compete with the contribution from ETH-monotonicity $\Delta E_{n,f}$. We will find $\Delta E_{mc,f}>>\Delta E_{mc,q}$ in higher-d CFTs when taking the limit $\bar{E}\to 0$. In other words, ETH-monotonicity dominates the quantum fluctuations.
We compare the contribution from ETH-monotonicity with the contribution from quantum fluctuations. We calculate the quantity $\Delta E_{mc,q}/\Delta E_{\beta}$ in the limit $\bar{E}\to 0$ or $r_{h}\to 0$.
@@
We have used $S(\bar{E})=4(d-1)\mu^{(d-1)/(d-2)}/\tilde{G}_{N}$ and $\bar{E}=\mu/\tilde{G}_{N}$. We can see that $\overline{\Delta E}_{mc}$ in (4.2) grows faster that the above quantity as $r_{h}\to 0$. So, the contribution from ETH-monotonicity dominates the contribution from the quantum fluctuations for small black hole microstates in higher dimensions. What is interesting about $\Delta E_{mc,q}$ is that it is still a positive quantity so it further reinforces the Kelvin statement of the second law. This is tied to the fact that small black holes in higher dimensions have negative heat capacity.
The behaviour of $\Delta E_{mc,q}/\Delta E_{\beta}$ for $2$-d holographic CFT is drastically different and will be considered in section 7. We will find that the quantum fluctuations numerically dominate ETH-monotonicity. Note that both their contributions are already exponentially small due to the $f(\bar{E})$ given by $\eta=3/2$. It agrees with the absence of curvature singularity in the BTZ black hole. Curvature invariants like the Kretschmann scalar is a fixed constant everywhere. We still find that $f(\bar{E})$ is monotonically increasing in the $2$-d holographic CFT. The quasiparticle peaks also do not appear in the limit $r_{h}\to 0$ because $r_{h}=0$ BTZ black hole is different from the $3$-D global AdS spacetime.
@@ 6 ETH-monotonicity in higher-d holographic CFT
This gives us the functional form in (4.1). This expression of $f(\bar{E})$ matches our numerical results, as one can see in Figure 2. Figure 2(a) is plots for $\Delta=\Delta_{+}=5/2,l=0$ of 4-d holographic CFT. Similarly, Figure 2(a) is plots for $\Delta=\Delta_{+}=3,l=0$ of 4-d holographic CFT. We have used $\lim_{\Delta\to 3}\;\sin(\pi\Delta)\,\Gamma(2-\Delta)=-\pi$.
@@
In principle, one can obtain the analytic expressions of $f(\bar{E})$ for all $l>0$ in the limit at $\mu\to 0$. One needs to take into account the instanton part of the Nekrasov-Shatashvili free energy. It is not necessary for $l=0$. The cancellation of the various divergent functions should also be performed with extreme care.
@@ 7 ETH-monotonicity in two-dimensional holographic CFT
The exponential factor dominates the power law dependence on $\beta$ as $\bar{E}\to 0$. Hence, the changes in energy $\Delta E_{\beta}$ and $\Delta E_{mc}$ vanish exponentially in this limit. Note that $f(\bar{E})$ is an increasing function of $\bar{E}$. So, $2$-d CFTs still satisfy R2 of ETH-monotonicity.
@@
We have used $S(\bar{E})=\pi\mu^{1/2}/(2G_{N})$ and $\bar{E}_{mc}=\mu/(8G_{N})$. It is now a negative quantity because small BTZ black holes still have positive heat capacity. On the other hand, $\overline{\Delta E}_{mc}$ is given by
@@ 8 Conclusions
In this work, we show that higher dimensional holographic CFTs possess ETH-monotonicity. So, smaller black hole microstates in higher dimensions have stronger contribution from ETH-monotonicity to the second law of thermodynamics. ETH-monotonicity also dominates quantum fluctuations in small black hole microstates in higher dimensions. The black hole singularity is constructed of microstates for which ETH-monotonicity starts competing with the entropic factor. We show that the relative extra gain in energy measures the curvature at the black hole horizon. Two-dimensional holographic CFT does not possess all features of ETH-monotonicity. The gain in energy vanishes exponentially with shrinking horizon radius. Quantum fluctuations also numerically dominate ETH-monotonicity in this case. These results are in agreement with the absence of curvature singularity in the BTZ black hole. So, we have
ETH-monotonicity = curvature at the horizon = inequivalence of ensembles.
We have quantified these equivalences by calculating the relative extra gain in energy (4.2) from ETH-monotonicity. It also holds true for pure states within a microcanonical energy window. We expect that ETH-monotonicity will persist even in the ultimate quantum theory of gravity. Because it is a property of many-body quantum chaotic system which becomes more prominent with decreasing system size. In a large black hole, ETH-monotonicity got washed out by the large entropy, and correspondingly, the curvature at the horizon is small.
@@
    Energy is measured in units of $1/R$ where $R$ is the AdS radius. We have taken $R=1$. $\bar{E}$, $M$ and $S$ are order $O(1/G_{N})$ quantities. The microcanonical energy window $\delta E$ is order $O(1/\sqrt{G_{N}})$. $\mu$, $\omega$, $r_{h}$, $\beta$ and $T$ are order $O(1)$ quantities in appropriate units.
@@
    The gain in energy from the entropic contribution $\Delta E_{\beta}$ is order $O(1)$, as expected for perturbations with light operators. The extra gain in energy from ETH-monotonicity $\Delta_{mc,f}$ and the quantum fluctuations $\Delta E_{mc,q}$ are order $O(G_{N})$. But we have shown that $\Delta E_{mc,q}$ is parametrically much smaller that $\Delta E_{mc,f}$ in small black holes in higher dimensions. So, ETH-monotonicity dominates the quantum fluctuations in this regime. Correction at the level of the microcanonical energy window $\delta E$ arises at order $O(G^{3/2}_{N})$, so it does not affect our analysis.
@@
    One may be wondering about the contribution of quantum fluctuations in the change in energy starting from a canonical ensemble (3.4). But we are not interested in this correction because its effect will be further suppressed by another factor of $G_{N}$. For example, in the expression of the relative extra gain in energy $\overline{\Delta E}_{mc}$ in (3.12), the finite size effect from $\Delta E_{\beta}$ will be order $O(G^{2}_{N})$.
@@
    When calculating $f(\bar{E})$ using (2.5) in small black holes, one should ideally use microcanonical ensemble. It will introduce $\omega$ in the $\bar{E}$-dependence just like the ETH-monotonicity contribution in $\Delta E_{mc}$. But again, the correction from this minute difference will be order $O(G^{2}_{N})$.
@@
The overarching idea is that inequivalence of ensembles manifesting as ETH-monotonicity is equivalent to high curvature at the horizon in small black holes. It is not unreasonable to expect that this idea would apply also to small black holes in asymptotically flat spacetime as well as asymptotically de-Sitter (dS) spacetime. After all, small black holes have the same behaviour in any of these spacetimes.
@@ Appendix A Energy gain starting from a microcanonical ensemble
We will show that the extra gain in energy due to ETH-monotonicity persists in case of a microcanonical ensemble. For this exercise, we use a toy density matrix constructed out of two eigenstates $|n_{1}\rangle$ and $|n_{2}\rangle$ with corresponding energy eigenvalues $E_{1}$ and $E_{2}$ respectively. Consider $E_{1}<E_{2}$. The density matrix is
@@
The energy expectation value is $\bar{E}_{mc}=(E_{1}+E_{2})/2$. $\delta E=E_{2}-E_{1}$ is the energy window of the microcanonical ensemble. It is of order $O(1/\sqrt{G_{N}})$ but parametrically much smaller than $\sqrt{G_{N}}\bar{E}_{mc}$. Let us have a closer look at the Taylor expansion of the entropic terms in (3.7) for the two eigenstates. For simplicity let us consider for a single frequency $\omega$.
$S(E_{1}+\omega)-S(E_{1}+\omega/2)$ $=$ $S(\bar{E}_{mc}-\frac{\delta E}{2}+\omega)-S(\bar{E}_{mc}-\frac{\delta E}{2}+\omega/2)$
$=$ $\frac{\beta\omega}{2}-\frac{\omega\delta E}{4}\,S^{\prime\prime}(\bar{E})|_{\bar{E}=\bar{E}_{mc}}+\frac{3\omega^{2}}{8}\,S^{\prime\prime}(\bar{E})|_{\bar{E}=\bar{E}_{mc}}+O(G^{3/2}_{N})$
$S(E_{2}+\omega)-S(E_{2}+\omega/2)$ $=$ $S(\bar{E}_{mc}+\frac{\delta E}{2}+\omega)-S(\bar{E}_{mc}+\frac{\delta E}{2}+\omega/2)$
$=$ $\frac{\beta\omega}{2}+\frac{\omega\delta E}{4}\,S^{\prime\prime}(\bar{E})|_{\bar{E}=\bar{E}_{mc}}+\frac{3\omega^{2}}{8}\,S^{\prime\prime}(\bar{E})|_{\bar{E}=\bar{E}_{mc}}+O(G^{3/2}_{N})\$
The correction at the level of the microcanonical energy window is $O(\sqrt{G_{N}})$ times bigger than the $\omega^{2}$ quantum fluctuations terms which are of the order of $O(G_{N})$. We can Taylor expand $e^{\pm\omega\delta ES^{\prime\prime}/2}$ for the two energy eigenstates. The opposing signs cancel these corrections at $O(\sqrt{G_{N}})$. The non-zero term with $\delta E$ starts at $O(G^{3/2}_{N})$ which is dominated by the quantum fluctuations and the ETH-monotonicity at $O(G^{1}_{N})$. The quantum fluctuations term coming from the other $S^{\prime\prime}$ terms has been treated separately in section 5.
Now let us consider the ETH-monotonocity form factor.
@@
$=$ $\omega|\tilde{\lambda}(\omega)|^{2}\,e^{\beta\omega/2}\left[f(\bar{E}_{mc}-\delta E+\omega/2,\omega)^{2}+f(\bar{E}_{mc}+\delta E+\omega/2,\omega)^{2}\right]\$
$\delta E\sim O(1/\sqrt{G_{N}})$ while $\bar{E}_{mc}\sim O(1/G_{N})$. So, Taylor expansion again cancels the $\delta E$ terms leaving the ETH-monotonicity form factor intact. The higher derivative corrections $(f^{\prime 2}+ff^{\prime\prime})\delta E^{2}$ appear to contribute at order $O(G_{N})$ but $\delta E$ is taken to be parametrically much smaller than $\sqrt{G_{N}}\bar{E}_{mc}$, even though of order $O(1/\sqrt{G_{N}})$. Say we take $\delta E=\epsilon\sqrt{G_{N}}\bar{E}_{mc}$, then this higher derivative term will be suppressed by $\epsilon^{2}$ factor as compared to the contribution from ETH-monotonicity $\Delta E_{mc,f}$. This is the story for our toy density matrix constructed with two eigenstates. Actually, this higher derivative term is order $O(G^{3/2}_{N})$ if we take a large number of eigenstates in the energy window. Let us say we take $M$ eigenstates, where $M$ is $O(1/\sqrt{G_{N}})$ number. This is the minimum number needed to pick up energy from perturbation with light operators. Because the energy gap between the eigenstates will be of $\delta E/M\sim O(1)$. So in the large $M$ limit, the higher derivative corrections will be
$\frac{(f^{\prime 2}+ff^{\prime\prime})\delta E^{2}}{M}\sum_{m=1}^{M}\frac{1}{m^{2}}=\frac{(f^{\prime 2}+ff^{\prime\prime})\delta E^{2}}{M}\frac{\pi^{2}}{6}\sim O(G^{3/2}_{N}).$ (A.2)
So finally, the extra gain in energy due to ETH-monotonicity survives and dominates even in a microcanonical ensemble.

### 25 Mar 2026 v4

@@ 5 Quantum fluctuations
The Taylor expansions of $S(E_{n}+\omega)$ and $S(E_{n}+\omega/2)$ in (3.6) have higher derivative terms $S^{\prime\prime}(\bar{E})$, $S^{\prime\prime\prime}(\bar{E})$, etc. $S^{\prime\prime}(\bar{E})$ gives a correction which is of the order $O(G^{1}_{N})$. $S^{\prime\prime\prime}(\bar{E})$ and higher derivative terms are suppressed by higher positive powers of $G_{N}$. So, we will concentrate on the contribution from $S^{\prime\prime}(\bar{E})=-\beta^{2}/C$, where $C$ is the total heat capacity Murthy:2019fgs . It is due to quantum fluctuations which mimic thermal fluctuations. So far, we have been ignoring this term. We expand the exponential in powers of $G_{N}$. The leading contribution to the change in energy due to the quantum fluctuations is given by
@@
It can compete with the contribution from ETH-monotonicity $\Delta E_{n,f}$. We compare the contribution from the quantum fluctuations with the contribution from ETH-monotonicity. We calculate the quantity $\Delta E_{mc,q}/\Delta E_{\beta}$ in the limit $\bar{E}\to 0$ or $r_{h}\to 0$.
@@
We have used $S(\bar{E})=4(d-1)\mu^{(d-1)/(d-2)}/\tilde{G}_{N}$ and $\bar{E}=\mu/\tilde{G}_{N}$. We can see that $\overline{\Delta E}_{mc}$ in (4.2) grows faster that the above quantity as $r_{h}\to 0$. So, the contribution from ETH-monotonicity dominates the contribution from the quantum fluctuations for small black hole microstates in higher dimensions. What is interesting about $\Delta E_{mc,q}$ is that it is still a positive quantity so it further reinforces the Kelvin statement of the second law. This is tied to the fact that small black holes in higher dimensions have negative specific heat.
The behaviour of $\Delta E_{mc,q}/\Delta E_{\beta}$ for $2$-d holographic CFT is drastically different and will be considered in section 7. We will find that the quantum fluctuations negate and numerically dominate ETH-monotonicity. Both their contributions are already exponentially small due to the $f(\bar{E})$ given by $\eta=3/2$.
@@ 6 ETH-monotonicity in higher-d holographic CFT
This gives us the functional form in (4.1). This expression of $f(\bar{E})$ matches our numerical results, as one can see in Figure 2. Figure 2(a) is plots for $\Delta=\Delta_{+}=5/2,l=0$ of 4-d holographic CFT. Similarly, Figure 2(b) is plots for $\Delta=\Delta_{+}=3,l=0$ of 4-d holographic CFT. We have used $\lim_{\Delta\to 3}\;\sin(\pi\Delta)\,\Gamma(2-\Delta)=-\pi$.
@@
In principle, one can obtain the analytic expressions of $f(\bar{E})$ for all $l>0$ in the limit at $\mu\to 0$. One needs to take into account the instanton part of the Nekrasov-Shatashvili free energy. It is not necessary for $l=0$. The cancellation of the various divergent functions should also be performed with extreme care. We will not pursue this objective in the present work.
@@ 7 ETH-monotonicity in two-dimensional holographic CFT
This is the exponential fall-off we have mentioned below (1.1). In contrast, it reduces to one in higher-d CFTs due to negative specific heat. The exponential factor dominates the power law dependence on $\beta$ as $\bar{E}\to 0$. Hence, the changes in energy $\Delta E_{\beta}$ and $\Delta E_{mc}$ vanish exponentially in this limit. Note that $f(\bar{E})$ is an increasing function of $\bar{E}$. So, $2$-d CFTs still satisfy R2 of ETH-monotonicity.
@@
We have used $S(\bar{E})=\pi\mu^{1/2}/(2G_{N})$ and $\bar{E}_{mc}=\mu/(8G_{N})$. It is now a negative quantity because small BTZ black holes still have positive specific heat. On the other hand, $\overline{\Delta E}_{mc}$ is given by
@@
The results of 2-d CFT are in agreement with the absence of curvature singularity in the BTZ black hole. Curvature invariants like the Kretschmann scalar is a fixed constant everywhere. We still find that $f(\bar{E})$ is monotonically increasing in the $2$-d holographic CFT. The quasiparticle peaks also do not appear in the limit $r_{h}\to 0$ because $r_{h}=0$ BTZ black hole is different from the $3$-D global AdS spacetime.
@@ 8 Conclusions
In this work, we study the the enveloping function $f(\bar{E},\omega)$ of the fluctuation term in the ETH statement (1.1) for holographic conformal field theories. We use the $\bar{E}$-dependence of this function to identify and study black hole microstates. In particular, we are interested in the microstates of small black holes which have high curvature at the horizon. We also set down a set of criteria for the $\bar{E}$-dependence which we called ETH-monotonicity. It reinforces the second law of thermodynamics over and above the entropic contribution in microcanonical ensembles of finite systems. It is a manifestation of inequivalence of ensembles. We show that higher dimensional holographic CFTs possess ETH-monotonicity. So, smaller black hole microstates in higher dimensions have stronger contribution from ETH-monotonicity to the second law of thermodynamics. ETH-monotonicity also dominates other quantum fluctuations in small black hole microstates in higher dimensions. The black hole singularity is constructed of microstates for which ETH-monotonicity starts competing with the entropic factor.
We show that the relative extra gain in energy measures the curvature at the black hole horizon. This is expected in other black hole geometries with negative specific heat and satisfying ETH-monotonicity. Two-dimensional holographic CFT does not possess all features of ETH-monotonicity. The gain in energy vanishes exponentially with shrinking horizon radius. Other quantum fluctuations also negate and numerically dominate ETH-monotonicity in this case. These results are in agreement with the absence of curvature singularity in the BTZ black hole. So, we conclude that
inequivalence of ensembles = ETH-monotonicity = curvature at the horizon
in higher dimensional holographic conformal field theories. We have studied all competing terms to come to this conclusion. We have quantified these relations by calculating the relative extra gain in energy (4.2) from ETH-monotonicity. It also holds true for pure states within a microcanonical energy window. We expect that ETH-monotonicity will persist even in the ultimate quantum theory of gravity. Because it is a property of many-body quantum chaotic system which becomes more prominent with decreasing system size. In a large black hole, ETH-monotonicity got washed out by the large entropy, and correspondingly, the curvature at the horizon is small.
@@
    Energy is measured in units of $1/R$ where $R$ is the AdS radius. We have taken $R=1$. $\bar{E}$, $M$ and $S$ are order $O(1/G_{N})$ quantities. The microcanonical energy window $\delta$ is order $O(1/\sqrt{G_{N}})$. $\mu$, $\omega$, $r_{h}$, $\beta$ and $T$ are order $O(1)$ quantities in appropriate units. The microcanonical energy fluctuation is relevant only for experimental measurements. We have full theoretical control at orders $O(1/G_{N})$, $O(1)$ and $O(G^{1}_{N})$. The most interesting physics is at order $O(G^{1}_{N})$.
@@
    The gain in energy from the entropic contribution $\Delta E_{\beta}$ is order $O(1)$, as expected for perturbations with light operators. But it is parametrically much smaller than one because we take $\tilde{\lambda}(\omega)$ to be very small so that perturbative calculation works.
@@
    The extra gain in energy from ETH-monotonicity $\Delta_{mc,f}$ and the quantum fluctuations $\Delta E_{mc,q}$ are order $O(G^{1}_{N})$. But we have shown that $\Delta E_{mc,q}$ is parametrically much smaller that $\Delta E_{mc,f}$ in small black holes in higher dimensions. So, ETH-monotonicity dominates the quantum fluctuations in this regime. Correction at the level of the microcanonical energy window $\delta$ are suppressed by $e^{-1/G_{N}}$, so it does not affect our analysis.
@@
    One may be wondering about the contribution of the quantum fluctuation in the change in energy starting from a canonical ensemble (3.4). But we are not interested in this correction because we have taken the expression as it is. It has been called the entropic contribution. If one is still interested in this correction for large black holes with stable canonical ensemble, it can be easily calculated. It will be further suppressed by another factor of $G_{N}$.
  5. 5.
    When calculating $f(\bar{E})$ using (2.5) in small black holes, one should ideally use microcanonical ensemble. It will introduce $\omega$ in the $\bar{E}$-dependence just like the ETH-monotonicity contribution in $\Delta E_{mc}$. But again, the overall correction from this minute difference will be order $O(G^{2}_{N})$.
@@
We believe that our technique of using $f(\bar{E})$ to identify and study black hole microstates will have wide applications in black hole physics and, in general, in quantum gravity. It is order $O(G^{1}_{N})$ effect but it can be tracked with precision. It is not unreasonable to expect that this idea would apply to black holes in asymptotically flat spacetime as well as small black holes in asymptotically de-Sitter spacetime. After all, small black holes have the same behaviour in any of these spacetimes. The only requirement is that the system is strongly chaotic so that ETH statement holds true. Finally, it is also very tempting to wonder if the Big Bang singularity were a very highly energetic large “white hole”, such as the one we defined in footnote 2.
@@ Appendix A Energy gain starting from a microcanonical ensemble
We will show that the extra gain in energy due to ETH-monotonicity persists in case of a microcanonical ensemble. For illustrative purpose, we use a toy density matrix constructed out of two eigenstates $|n_{1}\rangle$ and $|n_{2}\rangle$ with corresponding energy eigenvalues $E_{1}$ and $E_{2}$ respectively. Consider $E_{1}<E_{2}$. The density matrix is
@@
The energy expectation value is $\bar{E}_{mc}=(E_{1}+E_{2})/2$. $\delta=E_{2}-E_{1}$ is the energy window of the microcanonical ensemble. It is of order $O(1/\sqrt{G_{N}})$ but parametrically much smaller than $\sqrt{G_{N}}\bar{E}_{mc}$. Let us have a closer look at the Taylor expansion of the entropic terms in (3.6) for the two eigenstates. For simplicity let us consider for a single frequency $\omega$.
$S(E_{1}+\omega)-S(E_{1}+\omega/2)$
$=$ $S(\bar{E}_{mc}-\frac{\delta}{2}+\omega)-S(\bar{E}_{mc}-\frac{\delta}{2}+\omega/2)$
$=$ $\frac{\beta\omega}{2}-\frac{\omega\delta}{4}\,S^{\prime\prime}(\bar{E})|_{\bar{E}=\bar{E}_{mc}}+\frac{3\omega^{2}}{8}\,S^{\prime\prime}(\bar{E})|_{\bar{E}=\bar{E}_{mc}}+\frac{\omega\delta^{2}}{16}\,S^{\prime\prime\prime}(\bar{E})|_{\bar{E}=\bar{E}_{mc}}+O(G^{2}_{N})$
$S(E_{2}+\omega)-S(E_{2}+\omega/2)$
$=$ $S(\bar{E}_{mc}+\frac{\delta}{2}+\omega)-S(\bar{E}_{mc}+\frac{\delta}{2}+\omega/2)$
$=$ $\frac{\beta\omega}{2}+\frac{\omega\delta}{4}\,S^{\prime\prime}(\bar{E})|_{\bar{E}=\bar{E}_{mc}}+\frac{3\omega^{2}}{8}\,S^{\prime\prime}(\bar{E})|_{\bar{E}=\bar{E}_{mc}}+\frac{\omega\delta^{2}}{16}\,S^{\prime\prime\prime}(\bar{E})|_{\bar{E}=\bar{E}_{mc}}+O(G^{3/2}_{N})\$
The quantum fluctuations term $\omega^{2}S^{\prime\prime}$ has been treated separately in section 5. Taylor expansion of the exponentials of the remaining terms will kill all the odd $\delta$ power terms. It appears that terms like $(\omega\delta S^{\prime\prime})^{2}$ and $\omega\delta^{2}S^{\prime\prime\prime}$ are order $O(G^{1}_{N})$ quantities which could be a matter of concern. But all these terms with $\delta$ dependence will be suppressed by $e^{-1/G_{N}}$. This is because we take $M=O(e^{1/G_{N}})$ microstates in a generic microcanonical ensemble. So, the probability of each microstate is $1/M$. Any factor of $\delta/2$ should be then replaced by the difference of the energy eigenvalue from the microcanonical energy expectation value. For example, the final contribution from the first non-trivial term from the Taylor expansion of $e^{\omega\delta^{2}S^{\prime\prime\prime}/16}$ is
$\frac{\omega\delta^{2}S^{\prime\prime\prime}}{4M}\sum_{m=1}^{M}\frac{1}{m^{2}}=\frac{\omega\delta^{2}S^{\prime\prime\prime}}{M}\frac{\pi^{2}}{6}\sim O(e^{-1/G_{N}}).$ (A.2)
We have taken large $M$ limit because we are working with many-body excited states. Now let us consider the ETH-monotonicity form factor, again for the toy density matrix for easier illustration.
@@
$=$ $\omega|\tilde{\lambda}(\omega)|^{2}\,e^{\beta\omega/2}\left[f(\bar{E}_{mc}-\delta+\omega/2,\omega)^{2}+f(\bar{E}_{mc}+\delta+\omega/2,\omega)^{2}\right]\$
The $\delta$-dependent term starts from $(f^{\prime 2}+ff^{\prime\prime})\delta^{2}$. In a generic microcanonical ensemble, it is again suppressed by $e^{-1/G_{N}}$.